Let f (x) = 4x^2 + 2x - 8 and g(x) = 1 - x.
(a) Evaluate the function g - f for x = - 2. That is, find (g - f)( - 2).

(b) Evaluate the function fg for x = 2. That is, find (f g)(2).

(c) Find the difference function (f - g)(x) and simplify the results.


Answer by ikleyn(52781)

.

(a)  Evaluate each function f(x) and g(x) separately at x= -2; then  take the difference f(-2) - g(-2).


(b)  Evaluate each function f(x) and g(x) separately at x= -2; then  take the product  f(-2)*g(-2).


(c)  (f-g)(x) = by the definition = f(x) - g(x) = (4x^2 + 2x -8) - (1 - x) = 4x^2 + 3x -9.

Solved, answered and explained.